OpenStudy (anonymous):

How do you convert 85, 100 in/day to km/sec?

OpenStudy (theeric):

Do you know the general rule for converting with a conversion factor?

OpenStudy (anonymous):

not rally can be confusing

OpenStudy (theeric):

I know, it can be. But it all works because of the idea that you can multiply anything by \(1\) without changing the actual value. So, do you know how many seconds are in a day?

OpenStudy (anonymous):

86, 400 or 1440?

OpenStudy (theeric):

It is... I'll just do the calculation for myself. \(1\ day \times \dfrac{24\ hours}{day}\times\dfrac{60\ minutes}{hour}\times\dfrac{60\ seconds}{minute}\\1\ \cancel{day} \times \dfrac{24\ \cancel{hours}}{\cancel{day}}\times\dfrac{60\ \cancel{minutes}}{\cancel{hour}}\times\dfrac{60\ seconds}{\cancel{minute}}\\=1\times 24\times 60\times 60\ seconds=86,400\ seconds\) So 86,400 seconds.

OpenStudy (theeric):

Now, \(86,400\ seconds = 1\ year\). So, since they're equal, dividing one by the other will get you the amount of \(1\). Do you agree?

OpenStudy (anonymous):

Dividing 86,400 by 86, 400 equals 1, is that what you're talking about?

OpenStudy (theeric):

Oh! Before I meant "\(86,400\ seconds = 1\ day\)." I mixed up the day for year. That's true! And so is this: divide both sides by \(86,400\ seconds\) to get \(\dfrac{86,400\ seconds}{86,400\ seconds}=\dfrac{1\ day}{86,400\ seconds}\) which is the same as \(\dfrac{\cancel{86,400\ seconds}}{\cancel{86,400\ seconds}}=\dfrac{1\ day}{86,400\ seconds}\) \(1=\dfrac{1\ day}{86,400\ seconds}\)

OpenStudy (theeric):

Since \(1\ day\) and \(86,400\ seconds\) have the same value, one divided by the other will be equal to one. It's like the basic \(\dfrac{x}{x}=1\).

OpenStudy (anonymous):

ok, I see what you're saying

OpenStudy (theeric):

Okay! So, that value is called the conversion factor! Because, you can use it to make units cancel WITHOUT changing the VALUE! :) \(85,100\ \dfrac{in}{day}\times 1\\=85,100\ \dfrac{in}{day}\times \dfrac{1\ day}{86,400\ seconds}\\=85,100\ \dfrac{in}{\cancel{day}}\times \dfrac{1\ \cancel{day}}{86,400\ seconds}\\\approx .985\ \dfrac{in}{seconds}\) See? I mean, you still have to go from inches to kilometers, but that is half of the conversion!

OpenStudy (theeric):

Google says that \(39370.1\ seconds=1\ kilometer\). So, make a conversion factor!

OpenStudy (anonymous):

Ok, thanku a lot

OpenStudy (theeric):

If it's upside down, you'll just flip it. You want inches on the bottom, so that they cancel out with the inches on the top. Your very welcome!

OpenStudy (theeric):

You're*